how would you put 4x^2+5x=(3x-1)^2-3x^2 in the standard ax^2+bx+c=0 form
Answer by TimothyLamb(4379)

---
4x^2 + 5x = (3x - 1)^2 - 3x^2
7x^2 + 5x = (3x - 1)^2
7x^2 + 5x = (3x - 1)(3x - 1)
7x^2 + 5x = 9x^2 - 6x + 1
---
answer:
2x^2 - 11x + 1 = 0
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php